OpenStudy (steve816):

Water is poured into a container in the shape of a right circular cone with radius 4 feet and height 16 feet. Express the volume V of the water in the cone as a function of the height h of the water.

OpenStudy (campbell_st):

well if the height is h = 16 then the radius is r = 1/4 of the height so \[r = \frac{h}{4}\] so substitute these values into the formula \[V = \frac{1}{3} \pi r^2 h\] replace \[r ~~with~~\frac{h}{4}\] hope it helps

OpenStudy (campbell_st):

oops... forgot to say... simplify the equation

OpenStudy (steve816):

Thanks :)